# Ara-HOPE: Human-Centric Post-Editing Evaluation for Dialectal Arabic to Modern Standard Arabic Translation

### v1

@@ Abstract
Dialectal Arabic to Modern Standard Arabic (DA-MSA) translation is a challenging task in Machine Translation (MT) due to significant lexical, syntactic, and semantic divergences between Arabic dialects and MSA. Existing automatic evaluation metrics and general-purpose human evaluation frameworks struggle to capture dialect-specific MT errors, hindering progress in translation assessment. This paper introduces Ara-HOPE, a human-centric post-editing evaluation framework designed to systematically address these challenges. The framework includes a five-category error taxonomy and a decision-tree annotation protocol. Through comparative evaluation of three MT systems (Arabic-centric Jais, general-purpose GPT-3.5, and baseline NLLB-200), Ara-HOPE effectively highlights systematic performance differences between these systems. The results show that dialect-specific terminology and semantic preservation remain the most persistent challenges in DA-MSA translation. Ara-HOPE establishes a new framework for evaluating Dialectal Arabic MT quality and provides actionable guidance for improving dialect-aware MT systems.
@@
Abdullah Alabdullahthanks: The previous affiliation for Abdullah Alabdullah (where most of this research was completed) is the University of Manchester, Manchester, M13 9PL, The United Kingdom. School of Informatics University of Edinburgh The United Kingdom a.alabdullah@sms.ed.ac.uk          Lifeng Han Leids Uni. Medisch Centrum & LIACS Universiteit Leiden, NL l.han@lumc.nl          Chenghua Lin School of Computer Science University of Manchester The United Kingdom chenghua.lin@manchester.ac.uk
@@ 1 Introduction
Dialectal Arabic (DA) refers to the informal varieties of Arabic used in everyday communication, which vary significantly across regions. In contrast, Modern Standard Arabic (MSA) is the formal variety used in writing, education, and traditional media diab2010colaba.
Several key differences between DA and MSA affect MT: (i) Orthographic differences: Dialects do not follow standardized spelling rules and the same word may appear in different forms, making automatic text normalization difficult for NLP systems alhafni-etal-2024-exploiting. (ii) Morphological differences: Morphology is how words change form to express features such as tense or gender. While MSA has a rich, complex, and standardized morphology, spoken dialects often simplify these systems by omitting rules or using reduced forms KIRCHHOFF2006589. (iii) Lexical differences: Dialectal vocabulary often includes slang and idiomatic expressions that are typically absent in MSA. (iv) Syntactic differences: Syntax, sentence structure, and word order in many dialects differ from MSA 10.5555/1621774.1621784. (v) Code-switch: Speakers often mix DA and MSA, and sometimes even foreign words, within a single sentence. This adds more complexity when trying to build systems that automatically translate from DA to MSA hamed-etal-2025-survey. These challenges make DA-MSA translation a complex task requiring specialized approaches grounded in a sound understanding of common error types in DA-MSA translation systems.
In this paper we introduce a post-editing human evaluation framework that offers the granularity needed to identify systemic weaknesses in DA-MSA machine translation systems. Unlike general-purpose frameworks, our proposed framework (Ara-HOPE) specifically targets translation errors that result from DA-specific translation challenges.
@@ 2 Related Work
2.1 Advancements in DA translation
@@
Neural architectures have transformed DA-MSA translation by capturing contextual dependencies and complex syntactic and semantic relationships, effectively modeling dialectal variations and producing more fluent and accurate translations 9173dc9888094080b48ab529a68bbc63. The emergence of large parallel corpora has further advanced Neural Machine Translation by mitigating data scarcity and enabling training on dialectally diverse data. A key example is the MADAR corpus, which contains parallel translations from 25 Arabic city dialects Bouamor2018TheMA. Recent studies have compared neural architectures, including encoder–decoder models like NLLB-200 and decoder-only models like GPT-3 nllbteam2022languageleftbehindscaling; NEURIPS2020_1457c0d6. Decoder-only models have demonstrated better performance at preserving cultural context yakhni-chehab-2025-llms.
@@
The emergence of LLMs trained on large and diverse multilingual data, and further optimized through instruction tuning, has enabled models to follow natural-language instructions and generate appropriate outputs for a wide range of tasks, including machine translation, without costly task-specific fine-tuning NEURIPS2020_1457c0d6. This paradigm is known as In-Context Learning (ICL), where the desired task is specified directly in the prompt, and the model infers the mapping from the provided context. In zero-shot ICL, the model is instructed to translate from a source to a target language without providing any in-prompt examples. Zero-shot prompting has been particularly effective when parallel data is scarce and is widely used in DA-MSA shared tasks such as OSACT osact2024 and NADI abdulmageed2024nadi.
@@
While multilingual LLMs capture general linguistic features through large-scale multilingual pretraining, new Arabic-specialized LLMs like Jais improved the handling of dialectal nuances and cultural references sengupta2023jaisjaischatarabiccentricfoundation; mousi-etal-2025-aradice, producing more natural and contextually appropriate translations. Despite these advances, neural models continue to underperform on DA translation due to persistent challenges with dialectical nuances and culturally embedded expressions mousi-etal-2025-aradice.
2.2 DA translation evaluation
@@
Traditional automatic evaluation metrics such as BLEU and METEOR are limited for DA-MSA translation, as they rely on lexical overlap and perform poorly on morphologically rich languages like Arabic. Bouamor2014AHJ proposed AL-BLEU, an extension of BLEU that assigns partial credit for stem and morphological matches, yielding better correlation with human judgment than standard metrics. However, AL-BLEU remains a lexical-overlap metric and fails to capture semantic adequacy, particularly an issue for syntactically flexible languages like Arabic. While metrics like BLEU allow evaluation against multiple reference translations, these are challenging to produce. As a result, these metrics favor literal translations with high lexical overlap over culturally appropriate ones that better reflect human judgment, but differ lexically from the reference.
@@
In low-resource settings such as dialectal Arabic, neural evaluation metrics like BERTScore and COMET also face challenges. falcao2024comet showed that COMET’s performance on under-resourced languages is constrained by imbalanced training data. These metrics rely on pretrained models which typically lack sufficient dialectal Arabic training data, leading to lower-quality embeddings.
@@
Human evaluation frameworks such as the Multidimensional Quality Metrics (MQM) assess fine-grained translation errors like omissions and register mismatches lommel2014mqm; lommel-etal-2024-multi. However, MQM requires extensive annotator training kocmi-etal-2024-error. Moreover, DA translation requires a specialized human evaluation framework that captures the most impactful error types and minimizes subjectivity in assessing translation quality for this specific task, with only minimally trained annotators. To further advance the research in this direction, we will build our methodology on top of the HOPE metric gladkoff2022hope, a human-centric post-editing metric simplified from MQM, which has eight error types and error severity-level design.
@@ 3.1 Framework Development
Ara-HOPE Error Taxonomy Design: Developing a robust human evaluation framework requires clearly defined objectives. For DA-MSA translation, this requires an error taxonomy tailored to the specific challenges of this language pair. While the HOPE framework gladkoff2022hope offers a foundation for general post-editing translation assessment, it must be adapted to capture the specific linguistic phenomena of DA-MSA translation. Our proposed taxonomy is designed to evaluate translation quality, identify system weaknesses, and guide system improvements, while remaining usable by native DA speakers without requiring extensive annotation training.
@@
Design Principles: To ensure a theoretically sound and practical taxonomy, we developed guidelines grounded in best practices in translation quality assessment Han02072020; 10.1007/s10579-021-09537-5 and refined them through iterative feedback from focus groups with native DA speakers. The finalized guidelines comprise six core principles: (1) Identifiability: Errors must be detectable by minimally trained native dialect speakers to ensure consistent evaluation. (2) Distinguishability: Categories should have minimal conceptual overlap, with clear definitions to avoid confusion. (3) Actionability: Error classification must enable targeted system improvements through quantifiable error severity and clear mapping to issues in the translation system. (4) Comprehensiveness: The taxonomy must capture all major errors common in DA-MSA MT, including dialect-specific challenges. (5) Relevance: Categories must address issues unique to DA-MSA translation rather than pure general translation errors. (6) Usability: The taxonomy should remain manageable in size, balancing high-level categories with sufficient granularity.
@@
Taxonomy Structure: The Ara-HOPE taxonomy categorizes DA-MSA translation errors into three hierarchically structured types: (1) Fluency Error (FLU): Grammatical or linguistic errors in the MSA translation, independent of the DA source sentence. (2) Meaning Transfer Error: Failures to accurately preserve source meaning. This category includes Proper Name (PRN) errors, referring to incorrect translations of names of people, places, or organizations; Dialect-Specific Term (TRM) errors, which involve untranslated or mistranslated dialectal expressions that alter meaning; and General Semantic Mistranslation (GSMIS) errors, which includes omissions, additions, or other semantic changes. (3) Adaptation Error (ADP): Translations that are unnatural or contextually inappropriate in tone, style, or intent.
@@
GSMIS captures global meaning changes arising from DA’s contextual dependencies and includes all meaning changes not covered by PRN or TRM. This category was introduced after observing that LLMs often produce semantic distortions beyond proper name or dialect-specific term mistranslations.
@@
Decision Tree Implementation While a list format effectively outlined the taxonomy, it was later reformatted as a decision tree to reduce cognitive load for our minimally trained evaluators. The decision tree guides annotators through hierarchical error categories, as shown in Figure 5.
@@
This structure simplifies annotation by providing step-by-step guidance, helping evaluators understand the evaluation process, handle multiple error types systematically, and make consistent judgment at each level. The tree begins with the three primary categories (Fluency, Meaning Transfer, and Adaptation) and expands into more specific subcategories. Each node presents a yes/no question to minimize subjective judgments.
@@
Practical testing demonstrated Ara-HOPE’s effectiveness in addressing dialect-specific translation challenges and improving annotation consistency. The decision-tree structure also proved effective in guiding evaluators and enhancing usability.
@@ 3.2 Dataset Preparation
Our human evaluation experiment utilized 200 tweets from the Levantine development set of the Dial2MSA-Verified dataset khered2025dial2msa, a high-quality parallel corpus for DA-MSA translation in the social media domain. It extends the original Dial2MSA dataset mubarak2018dial2msa and applies automated corrections and human evaluation by native speakers to produce reliable MSA references.
@@
Our 200-example subset size exceeds the lower bound below which translation quality estimates become unreliable, addressing concerns that a very small sample may lead to unreliable error analysis gladkoff2022uncertainty.
@@ 3.3 Translation Systems Selection
The Selected Systems Are: Jais: A state-of-the-art Arabic-centric model trained with approximately one-third of its training tokens drawn from Arabic data, excelling at capturing DA-MSA nuances sengupta2023jaisjaischatarabiccentricfoundation. GPT-3.5: A general-purpose multilingual LLM. We will use this model to benchmark multilingual systems against an Arabic-specialized model. NLLB-200: A multilingual baseline MT system, providing contrast and revealing errors that more advanced systems may avoid nllbteam2022languageleftbehindscaling.
A zero-shot instruction setup was used to fairly evaluate each model’s baseline capabilities without fine-tuning or advanced prompting configuration. Jais and GPT-3.5 received consistent prompts (in Arabic for Jais and English for GPT-3.5) while NLLB-200 required no prompting due to its built-in MT task support.
@@ 4 Implementation
Annotators received initial training through a 25-minute pre-recorded video and supplementary materials. Training covered: (1) an Error Taxonomy Guide: providing detailed definitions of error types, alongside diverse examples of correct and incorrect DA-MSA translations, and (2) a Decision Tree Workflow: annotators first read the source and gold translation to establish the intended meaning, then highlighted errors by severity starting with the most critical, and finally used the decision tree (Figure 5) to classify errors systematically.

### v2

@@ Abstract
Dialectal Arabic to Modern Standard Arabic (DA-MSA) translation is a challenging task in Machine Translation (MT) due to significant lexical, syntactic, and semantic divergences between Arabic dialects and MSA. Existing automatic evaluation metrics and general-purpose human evaluation frameworks struggle to capture dialect-specific MT errors, hindering progress in translation assessment. This paper introduces Ara-HOPE, a human-centric post-editing evaluation framework designed to systematically address these challenges. The framework includes a five-category error taxonomy and a decision-tree annotation protocol. Through comparative evaluation of three MT systems (Arabic-centric Jais, general-purpose GPT-3.5, and baseline NLLB-200), Ara-HOPE effectively highlights systematic performance differences between these systems. Our results show that dialect-specific terminology and semantic preservation remain the most persistent challenges in DA-MSA translation. Ara-HOPE establishes a new framework for evaluating Dialectal Arabic MT quality and provides actionable guidance for improving dialect-aware MT systems. For reproducibility, we make the annotation files and related materials publicly available at https://github.com/abdullahalabdullah/Ara-HOPE
@@
Abdullah Alabdullahthanks: The previous affiliation for Abdullah Alabdullah (where most of this research was completed) is the University of Manchester, Manchester, M13 9PL, The United Kingdom.1, Lifeng Han2, Chenghua Lin3, 1School of Informatics, University of Edinburgh, The United Kingdom 2Leids Universitair Medisch Centrum & LIACS, Universiteit Leiden, NL 3School of Computer Science, University of Manchester, The United Kingdom Correspondence: a.alabdullah@sms.ed.ac.uk & l.han@lumc.nl,liacs.leidenuniv.nl
@@ 1 Introduction
This paper enhances Dialectal Arabic to Modern Standard Arabic (DA-MSA) translation quality assessment through a human-centric evaluation framework, addressing key gaps in current methods for under-resourced language pairs.
Dialectal Arabic (DA) refers to the informal varieties of Arabic used in everyday communication, which vary significantly across regions. In contrast, Modern Standard Arabic (MSA) is the formal variety used in writing, education, and traditional media Diab et al. (2010).
Unlike interlingual translation tasks (e.g. translating from English to Arabic), DA-to-MSA translation is a dialect normalization task that introduces additional challenges arising from divergences between Arabic dialects and MSA. Key differences between DA and MSA that affect MT include: (i) Orthographic differences: Dialects do not follow standardized spelling rules and the same word may appear in different forms, making automatic text normalization difficult for NLP systems Alhafni et al. (2024). (ii) Morphological differences: Morphology is how words change form to express features such as tense or gender. While MSA has a rich, complex, and standardized morphology, spoken dialects often simplify these systems by omitting rules or using reduced forms Kirchhoff et al. (2006). (iii) Lexical differences: Dialectal vocabulary often includes slang and idiomatic expressions that are typically absent in MSA Hadj Mohamed et al. (2023). (iv) Syntactic differences: Syntax, sentence structure, and word order in many dialects differ from MSA Biadsy et al. (2009). (v) Code-switch: Speakers often mix DA and MSA, and sometimes even foreign words, within a single sentence. This adds more complexity when trying to build systems that automatically translate from DA to MSA Hamed et al. (2025). These challenges make DA-MSA translation a complex task requiring specialized approaches grounded in a sound understanding of common error types in DA-MSA translation systems.
In this paper we introduce a post-editing human evaluation framework that offers the granularity needed to identify systemic weaknesses in DA-MSA machine translation systems. Unlike general-purpose frameworks, our proposed framework (Ara-HOPE) targets translation errors that result from DA-specific translation challenges.
@@ 2 Related Work
2.1 Advancements in Dialectal Arabic Translation
@@
Neural architectures have transformed DA-MSA translation by capturing contextual dependencies and complex syntactic and semantic relationships, effectively modeling dialectal variations and producing more fluent and accurate translations Baniata et al. (2018). The emergence of large parallel corpora has further advanced Neural Machine Translation by mitigating data scarcity and enabling training on dialectally diverse data. A key example is the MADAR corpus, which contains parallel translations from 25 Arabic city dialects Bouamor et al. (2018). Recent studies have compared neural architectures, including encoder–decoder models like NLLB-200 and decoder-only models like GPT-3 and GPT-4o Team et al. (2022); Brown et al. (2020); Alabdullah et al. (2025). Decoder-only models have demonstrated better performance at preserving cultural context Yakhni and Chehab (2025).
@@
The emergence of LLMs trained on large and diverse multilingual data, and further optimized through instruction tuning, has enabled models to follow natural-language instructions and generate appropriate outputs for a wide range of tasks, including machine translation, without costly task-specific fine-tuning Brown et al. (2020). This paradigm is known as In-Context Learning (ICL), where the desired task is specified directly in the prompt, and the model infers the mapping from the provided context. In zero-shot ICL, the model is instructed to translate from a source to a target language without providing any in-prompt examples. Zero-shot prompting has been particularly effective when parallel data is scarce and is widely used in DA-MSA shared tasks such as OSACT Atwany et al. (2024) and NADI Abdul-Mageed et al. (2024).
@@
While multilingual LLMs capture general linguistic features through large-scale multilingual pretraining, new Arabic-specialized LLMs like Jais improved the handling of dialectal nuances and cultural references Sengupta et al. (2023); Mousi et al. (2025), producing more natural and contextually appropriate translations. Despite these advances, neural models continue to underperform on DA translation due to persistent challenges with dialectical nuances and culturally embedded expressions Mousi et al. (2025); Alabdullah et al. (2025).
2.2 Dialectal Arabic Translation Evaluation
@@
Traditional automatic evaluation metrics such as BLEU Post (2018) and METEOR Banerjee and Lavie (2005) are limited for DA-MSA translation, as they rely on lexical overlap and perform poorly on morphologically rich languages like Arabic. Bouamor et al. (2014) proposed AL-BLEU, an extension of BLEU that assigns partial credit for stem and morphological matches, yielding better correlation with human judgment than standard metrics. However, AL-BLEU remains a lexical-overlap metric and fails to capture semantic adequacy, particularly an issue for syntactically flexible languages like Arabic. While metrics like BLEU allow evaluation against multiple reference translations, these are challenging to produce. As a result, these metrics favor literal translations with high lexical overlap over contextually appropriate ones that better reflect human judgment, but differ lexically from the reference.
@@
In low-resource settings such as dialectal Arabic, neural evaluation metrics like BERTScore Zhang et al. (2019) and COMET Rei et al. (2020) also face challenges. Falcão et al. (2024) showed that COMET’s performance on under-resourced languages is constrained by imbalanced training data. These metrics rely on pretrained models which typically lack sufficient dialectal Arabic training data, leading to lower-quality embeddings.
@@
Human evaluation frameworks such as the Multidimensional Quality Metrics (MQM) assess fine-grained translation errors like omissions and register mismatches Lommel et al. (2013, 2024). While MQM allows for more precise diagnostic error analysis due to its detailed error taxonomy, this comes at the cost of increased complexity in the annotation framework, requiring extensive annotator training, which can be costly and difficult to achieve for low-resource languages Kocmi et al. (2024). Moreover, DA translation requires a specialized human evaluation framework that captures the most impactful error types and minimizes subjectivity in assessing translation quality for this specific task, with only minimally trained annotators. To further advance research in this direction, we build our methodology on the HOPE metric Gladkoff and Han (2022). HOPE is a task-oriented evaluation framework designed to address the limitations of both automatic evaluation metrics (e.g. BLEU) and highly fine-grained but complex human evaluation frameworks such as MQM. Ara-HOPE is human-centric because it only incorporates eight evaluation criteria that capture the most critical and recurring errors in translation between the Levantine Arabic dialect and MSA, reflecting translation quality as perceived by native speakers. The assigned error scores correspond to the post-editing effort required to bring the translation to an acceptable quality.
@@ 3.1 Framework Development
Ara-HOPE Error Taxonomy Design: Developing a robust human evaluation framework requires clearly defined objectives. For DA-MSA translation, this requires an error taxonomy tailored to the specific challenges of this language pair. While the HOPE framework Gladkoff and Han (2022) offers a foundation for general post-editing translation assessment, we adapt the general error types in HOPE to capture the specific challenges of DA-MSA translation. To minimize subjective judgment between annotators, we also reduce the severity scoring range to three levels: from 0 (no errors) to 2 (major error). Our proposed taxonomy is designed to evaluate translation quality, identify system weaknesses, and guide system improvements, while remaining usable by native DA speakers without requiring extensive annotation training. To achieve this, we employ a direct quality estimation approach that evaluates predefined aspects of translation quality (e.g. fluency) at the segment level using discrete severity scores. This design makes our framework easier to train annotators on and faster to apply than MQM.
@@
Design Principles: To ensure a theoretically sound and practical taxonomy, we developed guidelines grounded in best practices in translation quality assessment Han (2020); Rivera-Trigueros (2022) and refined them through iterative feedback from focus groups with native DA speakers. The finalized guidelines comprise six core principles: (1) Identifiability: Errors must be detectable by minimally trained native dialect speakers to ensure consistent evaluation. (2) Distinguishability: Categories should have minimal conceptual overlap, with clear definitions to avoid confusion. (3) Actionability: Error classification must enable targeted system improvements through quantifiable error severity and clear mapping to issues in the translation system. (4) Comprehensiveness: The taxonomy must capture all major errors common in DA-MSA MT, including dialect-specific challenges. (5) Relevance: Categories must address issues unique to DA-MSA translation rather than pure general translation errors. (6) Usability: The taxonomy should remain manageable in size, balancing high-level categories with sufficient granularity.
@@
Taxonomy Structure: The Ara-HOPE taxonomy categorizes DA-MSA translation errors into three hierarchically structured classes: (1) Fluency Error (FLU): Grammatical or linguistic errors in the MSA translation, independent of the DA source sentence. (2) Meaning Transfer Error: Failures to accurately preserve source meaning. This category includes Proper Name (PRN) errors, referring to incorrect translations of names of people, places, or organizations; Dialect-Specific Term (TRM) errors, which involve untranslated or mistranslated dialectal expressions that alter meaning; and General Semantic Mistranslation (GSMIS) errors, which includes omissions, additions, or other semantic changes. (3) Adaptation Error (ADP): Translations that are unnatural or contextually inappropriate in tone, style, or intent.
@@
GSMIS captures global meaning changes arising from the model’s inability to handle DA’s contextual dependencies and includes all meaning changes not covered by PRN or TRM. This category was introduced after observing that LLMs often produce semantic distortions beyond proper name or dialect-specific term mistranslations.
@@
Decision Tree Implementation While a list format effectively outlined the taxonomy, it was later reformatted as a decision tree to reduce cognitive load for our minimally trained evaluators. The decision tree guides annotators through hierarchical error categories, as shown in Figure 5 in the Appendix. An English translated version of the tree is provided in Figure 6 in the Appendix.
@@
This structure simplifies annotation by providing step-by-step guidance, helping evaluators understand the evaluation process, handle multiple error types systematically, and make consistent judgment at each level. The tree begins with the three primary categories (Fluency, Meaning Transfer, and Adaptation) and expands into more granular subcategories. Each node presents a yes/no question to minimize subjective judgments. Since some error types (e.g. Dialect-Specific Term) can be challenging to distinguish from other categories, we provided annotators with a table of annotation guidelines with illustrative examples to clarify the distinctions between error types. The guidelines can be found in Figure 7 in the Appendix.
@@
Practical initial testing demonstrated Ara-HOPE’s effectiveness in addressing dialect-specific translation challenges and improving annotation consistency. The decision-tree structure also proved especially effective in guiding evaluators and enhancing usability.
@@ 3.2 Dataset Preparation
Our human evaluation experiment utilized 200 tweets from the Levantine development set of the Dial2MSA-Verified dataset Khered et al. (2025), a high-quality parallel corpus for DA-MSA translation in the social media domain. It extends the original Dial2MSA dataset Mubarak (2018) and applies automated corrections and human evaluation by native speakers to produce reliable MSA references.
@@
Our 200-example subset size exceeds the lower bound below which translation quality estimates become unreliable, addressing concerns that a very small sample may lead to unreliable error analysis Gladkoff et al. (2022).
@@ 3.3 Translation Systems Selection
The Selected Systems Are: Jais: A state-of-the-art Arabic-centric model trained with approximately one-third of its training tokens drawn from Arabic data, excelling at capturing DA-MSA nuances Sengupta et al. (2023). GPT-3.5: A general-purpose multilingual LLM. We will use this model to benchmark multilingual systems against an Arabic-specialized model. This model was introduced as an improvement over its predecessor, GPT-3 Brown et al. (2020). NLLB-200 3.3B: A multilingual baseline MT system, providing contrast and revealing errors that more advanced systems may avoid Team et al. (2022).
Each of the three systems was prompted to translate the same 200 examples. A zero-shot instruction setup was used to fairly evaluate each model’s baseline capabilities without fine-tuning or advanced prompting configuration. Jais and GPT-3.5 received consistent prompts (in Arabic for Jais and English for GPT-3.5) while NLLB-200 required no prompting due to its built-in MT task support.
@@ 4.1 Annotator Selection and Training
Annotators received initial training through a 25-minute pre-recorded video and supplementary materials. Training covered: (1) an Error Taxonomy Guide: providing detailed definitions of error types, alongside diverse examples of correct and incorrect DA-MSA translations, and (2) a Decision Tree Workflow: annotators first read the source and gold translation to establish the intended meaning, then highlighted errors by severity starting with the most critical, and finally used the decision tree (Figure 6 in the Appendix) to classify errors systematically.